### v2

# Radio observations point to a moderately relativistic outflow in the fast X-ray transient EP241021a

Abstract
Fast X-ray transients (FXRTs) are short-lived X-ray outbursts with diverse progenitor scenarios, including compact object mergers, stellar core-collapses and tidal disruption events. The Einstein Probe (EP) has enabled the rapid discovery and follow-up of dozens of FXRTs, revealing that while some of them overlap with traditional gamma-ray bursts (GRBs), a larger fraction of FXRTs have no associated gamma-ray counterpart down to deep limits. The origin of these gamma-ray dark FXRTs and their connection to the diverse landscape of stellar explosions remains an open question, which can be tackled through the study of their multi-wavelength counterparts and environment.
In this paper, we present long-term radio observations of the gamma-ray dark EP241021a, which exhibits sustained radio emission for over 100 days, placing it among the longest-lived radio afterglows. We detect signature of interstellar scintillation in early epochs, allowing us to constrain the angular size and Lorentz factor of the emitting region. Our observations point to an outflow that is at least mildly relativistic with $\Gamma\gtrsim 4$. Afterglow modeling favors a moderately relativistic ($\Gamma\approx$40) top-hat jet ($\theta_{c}\approx$0.16 rad) interacting with a low-density interstellar medium. The derived beaming-corrected kinetic energy, $E_{\mathrm{k}}\sim 1.6\times 10^{51}$ erg, and low radiative efficiency, $\eta_{\gamma}\lesssim$1%, are consistent with a standard relativistic explosion which did not produce bright gamma-rays. Alternatively, a highly-relativistic ($\Gamma\gtrsim$100) structured jet remains consistent with our observations if seen substantially off-axis ($\theta_{\rm obs}>\theta_{c}$). In the latter case, the initial X-ray flare detected by EP would be caused by the lower-$\Gamma$ ejecta from the lateral wings intercepting our line of sight rather than by traditional prompt-emission mechanisms within the jet core.
X-rays: bursts – radio continuum: transients – relativistic processes – stars: jets
1 Introduction
Fast X-ray transients (FXRTs) are sudden and intense bursts of X-ray emission occurring on timescales ranging from a few seconds to several hours (V. A. Arefiev et al., 2003; F. E. Bauer et al., 2017; J. Quirola-Vásquez et al., 2022). While FXRTs have been extensively investigated within the Galactic environment (A. Connors, 1988; D. M. Smith et al., 2006; V. Sguera et al., 2016; C. Ferrigno et al., 2019), their extragalactic counterparts remain poorly understood.
Extragalactic FXRTs typically exhibit peak luminosities spanning $10^{42}$ to $10^{50}$ erg s-1, placing them among the most luminous transient phenomena in the X-ray regime. Space-based X-ray observatories such as Swift (D. N. Burrows et al., 2005), Chandra (M. C. Weisskopf et al., 2002), XMM-Newton (F. Jansen et al., 2001), and eROSITA (P. Predehl et al., 2021) have played a critical role in their discovery and characterization (e.g., A. Glennie et al., 2015; Y. Q. Xue et al., 2019; D. Lin et al., 2020; S. Ai & B. Zhang, 2021; D. Lin et al., 2022). However, the long delay between the transient onset and its identification in archival searches often prevented placing meaningful constraints on their physical origin. A variety of progenitor scenarios, including compact object mergers, supernova shock break-outs, proto-magnetars, or tidal disruption events (TDEs) remain possible (Y. Q. Xue et al., 2019; G. Novara et al., 2020; D. Lin et al., 2022; J. Castaneda Jaimes et al., 2024). A notable case of a FXRT was reported by A. M. Soderberg et al. (2008), who discovered the X‐ray transient, XRF080109, in NGC 2770 at a distance of 27 Mpc. With a peak luminosity of  $\approx\,6\times 10^{43}\,\mathrm{erg\,s^{-1}}$, XRF080109 featured a 600-s X-ray flash followed by the emergence of supernova SN2008D, providing the first robust link between FXRTs and the core-collapse of massive stars (A. M. Soderberg et al., 2008; M. Modjaz et al., 2008; P. A. Mazzali et al., 2008). More recently, (F. E. Bauer et al., 2017) reported the discovery of the FXRT CDF-S XT1 in the Chandra Deep Field-South, characterized by a rapid rise of  $\sim\,100\,\mathrm{s}$, well described by a power-law decay with a slope of $\alpha\approx-1.53$, and a relatively hard spectrum with a photon index of 1.43. FXRT CDF-S XT1 was localized to a faint dwarf host galaxy at a redshift of $z=2.76$ (J. Quirola-Vásquez et al., 2024), suggesting a much more energetic explosion than XRF080109.
The advent of the Einstein Probe (EP; W. Yuan et al. 2022, 2025) has significantly advanced our understanding of FXRTs. By enabling the real-time discovery and localization of these events, EP has facilitated the identification of dozens of new FXRTs. Some of them, such as EP240219A (Y.-H. I. Yin et al., 2024), EP240315A (Y. Liu et al., 2025; R. Ricci et al., 2025; J. H. Gillanders et al., 2024), and EP240801a (S.-Q. Jiang et al., 2025), have been linked to classical gamma-ray bursts (GRBs; C. Kouveliotou et al., 1993; S. E. Woosley & J. S. Bloom, 2006). However, most FXRTs do not feature a gamma-ray counterpart, i.e., gamma-ray-dark FXRTs (e.g., EP240414a; H. Sun et al., 2024a). A subset of these gamma-ray dark FXRTs may originate from extremely distant GRBs, whose spectral peaks are redshifted into the X-ray band (J. Heise, 2003; V. D’Alessio et al., 2006); in other cases, such as EP240414A ($z=0.4$; H. Sun et al., 2024a; S. Srivastav et al., 2025; J. N. D. van Dalen et al., 2024), the events occur at relatively low redshifts yet still lack detectable gamma-ray counterparts, further challenging our understanding of their physical origins. A leading class of models attributes such gamma-ray dark FXRTs to relativistic outflows powered by a central engine. For example, off-axis jets, dirty fireballs, and choked jets associated to either GRBs or TDEs can reproduce the observed timescales and luminosities of FXRTs (J. E. Rhoads, 2003; B. Paczyński, 1998; C. D. Dermer, 1999; P. Mészáros & E. Waxman, 2001). Alternatively, models involving supernova shock breakouts, classical TDEs, or fast blue optical transients (FBOTs) may produce X-ray transients without invoking relativistic jets (A. M. Soderberg et al., 2008; T. W. S. Holoien et al., 2016; S. J. Prentice et al., 2018).
In this context, radio observations serve as a powerful diagnostic for distinguishing among these scenarios. Unlike optical emission, the radio band probes fast-moving ejecta from early to very late ($\gtrsim 1$ month) timescales, tracing long-lived synchrotron emission from relativistic electrons in the blast wave. These measurements provide valuable constraints on the energy budget and expansion dynamics of the outflow over extended periods (see e.g., R. Ricci et al., 2025; J. S. Bright et al., 2025). As radio waves propagate through the interstellar medium (ISM), they are subject to various propagation effects, including interstellar scintillation (ISS; see e.g., L. C. Lee & J. R. Jokipii 1975; J. W. Armstrong 1984).
This phenomenon has been extensively studied in compact radio sources such as pulsars and active galactic nuclei (AGN; B. J. Rickett, 1990; J. Goodman, 1997), and has also been observed in several GRB afterglows, (D. A. Frail et al., 1997, 2000; P. Chandra et al., 2008; J. Granot, 2008). ISS arises from small-scale electron density fluctuations in the ISM, which may induce rapid variations in signal intensity if the emitting source is sufficiently compact. Therefore, ISS provide important constraints on the angular size of the emitting region and, consequently, on the Lorentz factor of the outflow (D. S. Heeschen & B. J. Rickett, 1987; J. Dennett-Thorpe & A. G. de Bruyn, 2002; J. E. J. Lovell et al., 2008; B. J. Rickett et al., 1995), helping to determine whether the emission arises from a relativistic outflow.
This paper presents a study of EP241021a, a FXRT discovered by the Wide-field X-ray Telescope (WXT) onboard the EP on October 10, 2024 at $T_{0}=\text{05:07:56 UT}$, with a duration of approximately $100\,\mathrm{s}$ (J. W. Hu et al., 2024b). Notably, no gamma-ray counterpart was detected (D. Svinkin et al., 2024a) and the event was localized at a relatively low redshift of $z=0.75$ (see Fig. 1; G. Pugliese et al., 2024; W. Zheng et al., 2024; I. Pérez-Fournon et al., 2024). We carried out extensive radio follow-up of the afterglow, which remained detectable for over 100 days post-burst in the observer’s frame, placing it among the longest-lived radio afterglows to date (<< 4% of the events; P. Chandra et al., 2008).
[…]
2 Prompt Emission
Simultaneous X-ray and gamma-ray observations of the prompt emission are crucial to assess whether an event resembles a typical GRB spectrum or is intrinsically softer, suggesting an X-ray flash (XRF) or another non-standard origin (e.g., a supernova shock breakout). This diagnostic is especially important in our analysis of EP241021a and similar gamma-ray dark FXRTs, as it helps clarify whether the absence of gamma-ray emission arises from instrumental sensitivity limits or from genuinely distinct spectral characteristics.
Our dataset comprises 84 EP events observed up to March 21, 2025. Of these, 55 events have complete prompt emission coverage from the Fermi Gamma-ray Burst Monitor (GBM; C. Meegan et al., 2009) or were detections from Konus-Wind (R. L. Aptekar et al., 1995).
The remaining 29 events were either occulted by Earth or occurred while the detector was passing through the South Atlantic Anomaly (SAA), rendering prompt gamma-ray observations unavailable. In select cases, we extended the temporal search window by several minutes in an effort to recover partially observed events, thereby maximizing the number of FXRTs with usable prompt emission constraints.
Figure 2: Observed gamma-ray flux (10–1000 keV) versus X-ray flux (0.5–4 keV) for EP-discovered FXRTs. EP sources with joint gamma-ray detections (EP-GRBs) are shown by orange circles, and gamma-ray dark FXRTs by upward gray triangles (which mark $3\sigma$ upper limits to the gamma-ray emission, based on 8.192 s time-averaged fluxes from Fermi-GBM; (see Section 2). The X-ray fluxes represent the time-averaged, unabsorbed 0.5–4 keV fluxes derived from WXT observations; uncertainties are shown at the 90% confidence level. EP241021a is indicated by the downward red triangle. Dashed lines represent power-law spectra with photon indices $\Gamma=0.5$, $1.0$, and $1.5$.
Fig. 2 compares the (observer frame) gamma-ray (10–1000 keV) and X-ray (0.5–4 keV) fluxes for a sample of FXRTs, including EP241021a. For this study, $3\sigma$ upper limits on gamma-ray fluxes were derived using the Gamma-ray Targeted Search (GTS) [^1] https://github.com/USRA-STI/gamma-ray-targeted-search, built around the Gamma-ray Data Tools (A. Goldstein et al., 2024). We adopt a Band function [^2] The Band function is a phenomenological model of GRB spectra (D. Band et al., 1993). See Equation 1. defined as: $N(E)=\begin{cases}A\left(\dfrac{E}{100\,\mathrm{keV}}\right)^{\alpha}\exp\left(-\dfrac{E}{E_{0}}\right),&E<(\alpha-\beta)E_{0}\\ A\left(\dfrac{(\alpha-\beta)E_{0}}{100\,\mathrm{keV}}\right)^{\alpha-\beta}\exp(\beta-\alpha)\left(\dfrac{E}{100\,\mathrm{keV}}\right)^{\beta},&E\geq(\alpha-\beta)E_{0}\end{cases}$ (1) with the peak energy given by $E_{\rm peak}=(2+\alpha)E_{0}$, and we adopt typical GRB parameters $\alpha=-1.0$, $\beta=-2.3$, and $E_{\rm peak}=230~\mathrm{keV}$ (Normal template) to calculate the upper limits of energy fluxes in the case of non-detections. For events with detections, the time-averaged gamma-ray fluxes were taken directly from the values reported through GCN circulars (see Appendix A). The dashed lines in Fig. 2 indicate constant photon indices ($\Gamma=0.5$, $1.0$, and $1.5$), assuming a simple power-law spectrum. We observe that most gamma-ray bright FXRTs lie at $\Gamma\sim 1.0$, consistent with the region typically occupied by standard GRBs (S. Poolakkil et al., 2021).
For EP241021a, we used the X-ray flux (0.5–4 keV) estimated by J. W. Hu et al. (2024b) of $(3.3^{+1.5}_{-1.7})\times 10^{-10}~\mathrm{erg~cm^{-2}~s^{-1}}$. Using the targeted search, we place a $3\sigma$ upper limit on the gamma-ray flux of EP241021a as $3.2\times 10^{-8}~\mathrm{erg~cm^{-2}~s^{-1}}$ for a timescale of 8.192 s. At a redshift of $z=0.75$, this corresponds to a limit on the isotropic-equivalent gamma-ray energy of $E_{\gamma,\rm iso}\lesssim 7.2\times 10^{50}$ erg, consistent with the value inferred by M. Busmann et al. (2025) and at the lower end of the distribution for cosmological GRBs, although still consistent with it.
A small subset of EP events fall below the $\Gamma=1.5$ line, which lies at the higher end of the X-ray flux distribution. These events are particularly intriguing, as they are bright in X-rays yet lack detectable gamma-ray emission. This suggests that their gamma-ray faintness is not due to instrumental biases and may instead reflect unusually softer spectra. These are the best candidates to search for very high-redshift bursts, off-axis and choked jets, as well as exotic types of explosions.
The two classes of gamma-ray bright and gamma-ray dark transients mostly overlap in the region of $\Gamma=1.0-1.5$, where EP241021a also lies. Therefore, based solely on its high-energy properties, it is not possible to robustly identify EP241021a as either a GRB-like explosion or a peculiar X-ray rich transient.
In this context, the X-ray to gamma-ray flux ratio of EP241021a remains consistent with the broader EP-GRB population (Fig. 2). This indicates that prompt emission properties alone may often be insufficient to unambiguously determine the nature of FXRTs.
3 Afterglow Observations
3.1 Australia Telescope Compact Array
Australia Telescope Compact Array (ATCA) observations were conducted over 11 epochs between October 28, 2024, and January 25, 2025, covering frequencies from 2.1 GHz to 16.7 GHz. The bandwidth in all the observations was 2 GHz in the Compact Array Broadband Backend 1 MHz per frequency channel mode. The primary and bandpass calibrator was 1934-638 for all observations at 2.1, 5.5 and 9 GHz. At 16.7 GHz, 1921–293 served as the bandpass calibrator observations, while 1934-638 remained the primary calibrator. The phase calibrator 0146+056 was used to correct complex gains. The 21.2 GHz data were taken at the same time of the 16.7 GHz ones but the 21.2 GHz flux densities are not presented in this work because high system temperature values and large flux density scale corrections from the primary calibrator resulted in unreliable flux density values for the phase calibrator and the target source. Data reduction was performed using the Miriad (R. J. Sault et al., 1995) software package, which was used to flag Radio Frequency Interferences (RFIs), calibrate the complex gains, bootstrap the flux density scale, and image the fringe visibilities. A null robustness parameter value was used as a good trade-off between uniform and natural weighting in the imaging strategy. Final cleaned radio maps were inspected by Karma kvis (KARMA kvis, 1995) and ds9 (Smithsonian Astrophysical Observatory, 2000) visualisation softwares. When a detection was found the peak flux was extracted, otherwise, a 3-$\sigma$ upper limit was derived from the Root Mean Square (RMS) noise in the restored maps. The uncertainties on the flux densities were estimated by summing in quadrature the map RMS noise and 5% of the peak flux density, accounting for residual complex gain calibration errors.
Table 1: Summary of EP241021a: ATCA Flux Density Measurements at 2.1, 5.5, 9, and 17 GHz. The columns are as follows: Epoch, Start Time (UT), $T-T_{0}$ (days) representing the time elapsed since the burst detection, and $S_{\nu}$ representing the peak flux density at each observing frequency.
[…]
1 2024-10-29T10:43:25 8.3 5.5 356 $\pm$ 18
9 413 $\pm$ 21
2 2024-11-08T8:52:55 18.3 5.5 434 $\pm$ 22
9 543 $\pm$ 27
17 366 $\pm$ 35
3 2024-11-11T8:34:05 21.3 2.1 339 $\pm$ 31
4 2024-11-15T7:36:25 25.3 5.5 802 $\pm$ 42
9 596 $\pm$ 31
17 347 $\pm$ 41
5 2024-11-21T7:30:05 32.3 2.1 <117<117
5.5 937 $\pm$ 52
9 987 $\pm$ 54
6 2024-11-24T7:01:$0$ 34.1 17 450 $\pm$ 46
7 2024-12-07T8:29:45 47.3 5.5 620 $\pm$ 45
9 565 $\pm$ 43
17 361 $\pm$ 50
8 2024-12-08T8:06:35 48.3 2.1 761 $\pm$ 228
9 2024-12-23T7:01:25 63.2 2.1 403 $\pm$ 55
5.5 875 $\pm$ 49
9 870 $\pm$ 43
10 2024-12-30T5:36:55 70.1 2.1 395 $\pm$ 118
5.5 492 $\pm$ 34
9 366 $\pm$ 25
17 180 $\pm$ 54
11 2025-01-25T5:14:35 96.0 2.1 250 $\pm$ 60
5.5 428 $\pm$ 30
9 335 $\pm$ 24
Table 2: Summary of EP241021a: e-MERLIN Flux Density Measurements at 5 GHz. The columns are as follows: $T-T_{0}$ representing the time elapsed since the burst detection, and $S_{\nu}$ representing the peak flux density at the corresponding frequency.
[…]
  1   48.8   5   355 $\pm$ 34
  2   74.6   5   454 $\pm$ 41
  3   110.7   5   206 $\pm$ 35
  4   136.5   5   118 $\pm$ 27
3.2 e-MERLIN
EP241021a was observed with enhanced Multi Element Remotely Linked Interferometer Network (e-MERLIN) in four epochs at the centre frequency of 5 GHz with a total bandwidth of 512 MHz divided in 4 sub-bands of 128 MHz each. The data were correlated by the e-MERLIN software correlator. The resulting fringe visibilities were calibrated in the CASA software v.5.8 using the e-MERLIN CASA pipeline (e-MERLIN pipeline, 2019; J. Moldon, 2021) and imaged in CASA v.5.5.0 using a Briggs weighting scheme and a robustness value of 0.5. The target was observed together with 1331+3030 (alias 3C286), 1407+2827, 0154+0823 as the flux calibrator , the pointing calibrator and the bandpass calibrator, respectively. The target was fit with a 2D Gaussian in CASA and the peak flux was quoted in the flux density measurement, being it point-like in the $100\times 40$ mas final cleaned maps. The e-MERLIN position obtained by 2D Gaussian fitting to the cleaned map of EP241021a is RA = 01h 55m 23.4325s, Dec = +05 56 17.774′′ (J2000) with a positional $1\sigma$ uncertainity of (3, 9) mas in RA and DEC, respectively. Flux uncertainties were estimated as detailed in Table 1. Detections and the earlier flux upper limit obtained by this work are shown in Table 2.
3.3 Very Large Array
[…]
3.4 X-ray Observations
Observations of EP241021a were conducted with the Follow-up X-ray Telescope (FXT) onboard Einstein Probe, starting $\approx$36.5 hours after the EP-WXT trigger (Y. Wang et al., 2024a). Follow-up observations from the Swift X-ray Telescope (XRT) were reported by M. Busmann et al. (2025), revealing a long-lasting X-ray counterpart. To determine whether the long-lived nature of the X-ray emission was possibly due to contamination from a nearby source, we analyzed public data from the Chandra X-ray Observatory contained in the Chandra Data Collection https://doi.org/10.25574/cdc.415 (catalog DOI: 10.25574/cdc.415), which observed EP241021a approximately 14 days after trigger (ObsID: 30598; PI: Jonker) for a total exposure of 9.9 ks. A single X-ray source is detected within the XRT localization at coordinates RA = 01h 55m 23.3984s, Dec = +05 56 17.498′′ (RA = 28.847, Dec = +5.938) with $90\%$ positional uncertainty of $0.52^{\prime\prime}$. Source counts were extracted using the CIAO software package (v4.16) [^3] https://cxc.harvard.edu/ciao/, employing a circular aperture with a radius of 1.1 arcseconds, corresponding to approximately 90% encircled energy of the Chandra point spread function (PSF). Aperture photometry was performed in the 0.5–8.0 keV band. The resulting spectrum was modeled using an absorbed power-law model (tbabs*powerlaw) in XSPEC, with the hydrogen column density ($N_{\mathrm{H}}$) fixed to the Galactic value reported in J. W. Hu et al. (2024b). We find a best-fit photon index of $\Gamma=1.3\pm 0.7$, and an unabsorbed model flux of $(6.1^{+1.4}_{-1.2})\times 10^{-14}$ erg cm-2 s-1 in the 0.3–10 keV band. This value is consistent with the nearly simultaneous XRT measurements (M. Busmann et al., 2025) and confirms that the late-time flattening is an intrinsic property of the source.
Figure 3: ATCA and e-MERLIN radio observations of EP241021a. Flux densities are listed in Table 1 and 2. The inset panel displays the intra-observation light curve from the first epoch at 5.5 GHz, segmented into 40-min intervals to highlight short-term variability.
4 Constraints from radio observations
As mentioned in the Section 1, afterglows from compact radio sources are subject to ISS, which can produce significant flux modulations due to distorted wavefronts (M. A. Walker, 1998). Specifically, such flux variations can arise from diffractive and refractive scintillation, each sensitive to different angular scales and frequency bandwidths (R. Narayan, 1992; J. Goodman, 1997). By identifying the epochs when diffractive scintillation ceases and refractive scintillation becomes suppressed, we can place lower and upper bounds, respectively, on the angular size of the source. These size constraints, when combined with the angular diameter distance to the source, allow us to infer limits on the Lorentz factor of the outflow, as demonstrated in previous studies of GRBs and other relativistic transients (e.g., D. A. Frail et al., 2000; S. B. Cenko et al., 2008; L.-L. Zhang et al., 2023). Motivated by this, in Section 4.1, we present evidence for ISS during the radio monitoring of EP241021a. In Section 4.2, we use the ISS behavior to constrain the source’s angular size and Lorentz factor. We calculate the brightness temperature in Section 4.3 using early-time radio data.
4.1 Evidence for ISS
The strength and timescale of ISS effects depend on both the observing frequency and the angular size of the source (e.g., Y. Liu et al., 2022a). ISS manifests on two distinct timescales: diffractive ISS (DISS), which induces large-amplitude and rapid fluctuations over minutes due to small-scale interference patterns; and refractive ISS (RISS), which causes slower variability over weeks to months as the signal traverses larger-scale density structures in the scattering screen (M. A. Walker, 1998; W. Lewandowski et al., 2011; J. Granot & A. J. van der Horst, 2014; Y. Liu et al., 2022b). These scintillation effects can produce substantial variations in both flux and spectral shape, especially during the early stages of an afterglow. Initially, when the GRB jet is highly-collimated, DISS dominates the variability; as the emitting region expands, DISS becomes suppressed, and RISS emerges as the dominant contributor (D. A. Frail et al., 1997). Consequently, different radio facilities, operating at various frequencies and timescales, might detect variability on distinct timescales.
[…]
Assuming that scattering occurs in a thin screen along the line of sight, the screen is located at a distance $D_{\text{scr}}\sim(h_{z}/2)\sin^{-1}|b|\approx 0.6$ kpc from the observer (R. J. Reynolds, 1989; T. J. Galama et al., 2003), where $h_{z}$ is the scale height of the ionised gas layer ($\sim 1$ kpc) and b=53.47b=-53.47 is the Galactic latitude in degrees. We use the NE2001 model of the Galactic distribution of free electrons (J. M. Cordes & T. J. W. Lazio, 2002; M. A. Walker, 1998) to derive the scattering measure (SM) of $1.3\times 10^{-4}\,\text{kpc}\,\text{m}^{-20/3}$ and a transition frequency of $\nu_{0}\sim 6$ GHz. Below this frequency, density irregularities in the ISM cause strong phase perturbations leading to significant scattering, whereas above this frequency, scattering effects become weaker.
The decorrelation bandwidth ($\Delta\nu_{\rm dc}$) for diffractive scintillation represents the frequency range over which the intensity variations caused by diffraction remain correlated (J. Goodman, 1997):
$\Delta\nu_{\text{dc}}=2.2\left(\frac{\nu}{5.5\text{ GHz}}\right)^{22/5}\left(\frac{D_{\text{scr}}}{\text{kpc}}\right)^{-1}\left(\frac{\rm SM}{10^{-4}\text{ m}^{-20/3}\text{ kpc}}\right)^{-6/5}\text{ GHz}\,.$ (2)
Since $\Delta\nu_{\text{obs}}\sim\Delta\nu_{\text{dc}}$ = 2.5 GHz, diffractive scintillation is expected to have a significant influence at $\approx$5.5 GHz and be progressively suppressed at lower frequencies. Moreover, the 2.1 GHz frequency data have a lower signal-to-noise ratio which precludes a time-variability analysis.
4.2 Angular size measurements and Lorentz factor
The angular scales can be expressed in terms of the angular size of the first Fresnel zone evaluated at the transition frequency $\nu_{0}$. The characteristic angular size is given by:
$\theta_{F_{0}}=2.1\times 10^{4}\,{\rm SM}^{0.6}\,\nu_{0}^{-2.2}\,\mu\text{as}\,$, (3)
[…]
In the case of refractive scintillation, the scattering disk is much larger than the first Fresnel zone, and this results in $\theta_{r}=\theta_{\mathrm{F0}}(\nu/\nu_{0})^{-11/5}=2.3\,\mu\mathrm{as}$} at 5.5 GHz (J. Granot & A. J. van der Horst, 2014). Large amplitude fluctuations are expected at 5.5 GHz, with modulation index $m_{r}=(\nu/\nu_{0})^{17/30}\sim 0.95$ (J. Granot & A. J. van der Horst, 2014), and at 2.1 GHz with $m_{r}\sim 0.6$. This effect is observed up to $\approx 25$ d at 5.5 GHz and up to $\approx 32$ d at 2.1 GHz. It explains the large discrepancy between the measured fluxes at different frequencies, but is likely quenched afterwards.
By imposing $\theta_{s}\lesssim\theta_{r}$ at 25 d (5.5 GHz), we derive a rather tight upper limit on the physical size of the source $D_{s}\lesssim\theta_{r}D_{A}\approx 6\times 10^{16}$ cm. Following the steps outlined above, we derive an upper limit on the Lorentz factor of $\Gamma\lesssim$4.3 at 1.5 d (observer’s frame).
Therefore, the observed variability of the radio counterpart consistently points to an outflow that is at least mildly relativistic, with Lorentz factor $\Gamma\approx$4. Unfortunately, without a clear identification of the deceleration time (e.g., E. Molinari et al., 2007), we cannot exclude the presence of an ultrarelativistic outflow at earlier times.
Figure 4: Variability timescale on October 29, 2024 with data sampled at 40-minute intervals.
4.2.1 Variability function
The above estimates depend on the adopted ISM model, in particular the scattering measure. An alternative way to estimate this quantity directly from the data is by examining the variability timescale using intensity structure functions (J. H. Simonetti et al., 1985; P. A. Hughes et al., 1992; A. Ciaramella et al., 2004; P. Chandra et al., 2008),
$V=\frac{1}{N-1}\sum_{i=1}^{N-1}\left[F(t_{i}+\Delta t)-F(t_{i})\right]^{2}\,$, (4)
where $F(t)$ is the flux density at time $t$, $\Delta t$ is the fixed time interval between consecutive measurements, and $N$ is the total number of observations. This variability metric quantifies the mean squared deviation of the flux measurements from the reference value, serving as an indicator of the temporal variability induced by the scintillation. We tentatively identify a break with 20–30% accuracy at $\Delta T\approx 2\times 10^{4}$ s (see Fig. 4), which places a lower limit to the true timescale of diffractive scintillation $t_{\rm diff}$. For a relative transverse velocity $v_{\text{rel}}\sim 30$ km s-1, we derive a scattering measure:
${\rm SM}\lesssim 0.4\left(\frac{\nu}{5.5\text{ GHz}}\right)^{2}\left(\frac{t_{\text{diff}}}{20,000\text{ s}}\right)^{-5/3}\left(\frac{v_{\text{\rm rel}}}{30\text{ km s}^{-1}}\right)^{-5/3}\\ \times 10^{-4}\,\text{m}^{-20/3}\text{ kpc\,,}$ (5)
[…]
4.3 Brightness Temperature
The brightness temperature $T_{b}$ provides an estimate of the intensity of radiation emitted by a source, assuming it behaves like a blackbody. Its value can be derived from early-time radio observations to aid constraining the relativistic nature of the outflow (G. E. Anderson et al., 2014).
For EP241021a, we use the first ATCA detection at $\sim$8 days for calculating $T_{b}$ as (M. S. Longair, 2011):
$T_{b}=1.153\times 10^{-8}\,D_{L}^{2}S_{\nu}\nu^{-2}t^{-2}(1+z)^{-1}\approx 3.3\times 10^{13}K,$ (6)
where $S_{\nu}$ is the flux in Jy, $\nu$ is the frequency in Hz, $t$ is the time since the trigger in seconds, $z$ is the redshift. Since the maximum possible emitting region size is constrained by the speed of light ($ct$), comparing the observed $T_{b}$ with the inverse-Compton limit ($T_{B}\approx 10^{12}$ K) allows us to estimate the minimum bulk Lorentz factor ($\Gamma$) of the emitting material. This relationship is given by (T. J. Galama et al., 1999):
$\Gamma\gtrsim\left(\frac{T_{b}}{T_{B}}\right)^{1/3}\,\approx 3.$ (7)
Extrapolating this value back in time, using the relation $\Gamma\propto t^{-3/8}$ as discussed in the previous section, we derive a lower limit on the Lorentz factor $\Gamma\gtrsim$5.5 at 1.5 d. This is slightly larger than the constraints from ISS, however, considering the uncertainties in the models and underlying assumptions, both these results appear in agreement and consistently point to an outflow that is at least mildly relativistic.
5 Multi-wavelength Afterglow Modeling
In the previous sections, we demonstrated that radio observations point to an expanding outflow with at least mildly relativistic velocity. The interaction of this fireball with the surrounding medium is known to produce broadband synchrotron radiation (R. Sari et al., 1998; P. Mészáros et al., 1999; J. Granot & R. Sari, 2002), known as afterglow, whose brightness and temporal evolution depend, among other factors, on the outflow structure, its geometry, and the observer’s viewing angle (P. Rossi et al., 1993; R. Sari et al., 1999; E. Troja et al., 2019; G. Ryan et al., 2020, 2024). This afterglow emission is generally described with a power-law form: $F_{\nu}\propto\nu^{-\beta}t^{-\alpha}$, where $\alpha$ and $\beta$ obey to standard closure relations (R. Sari et al., 1999; B. Zhang & P. Mészáros, 2004). We use this empirical model to derive basic physical constraints from the afterglow observations.
The evolution of the radio spectral index, shown in Fig. 5, reveals the following:
Figure 5: Spectral index evolution between 5.5–9 GHz (red circles) and 9–17 GHz (blue squares) over time since the EP trigger. The dashed line indicates $\beta=0.75$.
  • the spectral index between 9 and 17 GHz remains constant around the value $\beta_{2}\sim 0.75$. This is consistent with the synchrotron regime $\nu_{m}<\nu_{r}<\nu_{c}$ and implies a characteristic synchrotron frequency $\nu_{m}\lesssim 9$ GHz after $T_{0}+18$ days, and a spectral index $p=2\beta_{2}+1\approx 2.5$ of the shock-accelerated electrons.
  • the spectral index between 5.5 and 9 GHz is seen to vary. At early times, its value might be affected by ISS, which would explain its rapid variations. However, it displays a consistent trend of increasing from $\beta_{1}\sim-1/3$ to $\beta_{2}\sim 0.75$. This is consistent with the characteristic synchrotron frequency $\nu_{m}$ crossing this frequency range and progressively moving below 5.5 GHz after $T_{0}+70$ days, as expected for an external forward shock expanding into a uniform medium.
The observed behavior is broadly consistent with expectations of an external forward shock expanding into a uniform medium, for which $\nu_{m}\propto t^{-3/2}$ and $F_{m}\approx$ constant, and rules out any major contribution from reverse shock.
We use the best-sampled radio light curve at 5.5 GHz, including our 6 GHz observations, to characterize the afterglow temporal evolution. A smoothly broken power-law fit (R. Ricci et al., 2025) yields a rising slope of $\alpha_{1}=0.7\pm 0.08$ (68% confidence level), peaking at $\sim$40 d (rest frame), followed by a decay with slope $\alpha_{2}=1.9\pm 0.1$ (68% c.l.).
The early rise, although likely affected by ISS, is consistent with the expected behavior for $\nu_{r}\lesssim\nu_{m}$, as also suggested by the spectral index $\beta_{1}\approx$-1/3. The steep late-time decay helps further constraining models as it disfavors quasi-spherical outflows, such as mildly relativistic cocoon or choked jet models (e.g., E. Troja et al., 2019), and points to a collimated outflow.
Our empirical analysis shows that the radio emission in the range 2-17 GHz is broadly consistent with the evolution of optically thin synchrotron radiation from a population of shock-accelerated electrons. This allows us to place an upper limit on the self-absorption frequency $\nu_{sa}\lesssim$5 GHz, which in turn provides an independent constraint on the bulk Lorentz factor, $\Gamma\approx$4 (R. Barniol Duran et al., 2013). In contrast, the optical and near-infrared counterpart of EP241021a display a much more complex behavior, with a prominent bump at $\approx$7 d followed by a shallow long-lasting decay (M. Busmann et al., 2025). Non-standard evolution is also observed at X-ray energies (Sec. 3.4). Therefore, we exclude the late-time ($\gtrsim$7 d) optical and X-ray data from our modeling and test whether a basic afterglow model can consistently describe the radio dataset and the early optical and X-ray observations.
Figure 6: Broadband modeling results. Best-fit afterglow light curves for an on-axis jet (model Top-Hat) compared with observations at multiple frequencies. The X-ray (Swift/XRT; 1 keV) and optical and near-infrared observations in the r, z and J filters are from M. Busmann et al. (2025). Radio observations at 15.5 GHz from AMI-LA (F. Carotenuto et al., 2024). Downward arrows indicate $3\sigma$ upper limits. Empty symbols show data not used for model fitting. A scaling factor has been applied to different frequencies for plotting purposes.
Figure 7: Broadband modeling results. Corner plot showing the posterior distributions of the fitted model parameters using a uniform (top-hat) jet model.
To derive physical constraints on the explosion, we model the afterglow using the open-source Python package Afterglowpy v. 0.8.0 (G. Ryan et al., 2020, 2024), which implements the single-shell approximation (H. van Eerten et al., 2010; H. van Eerten, 2018). In this framework, the ejecta and forward shock are treated as a single radially uniform fluid element. The model incorporates prescriptions for jet spreading and uses an equation of state that smoothly transitions between the ultra-relativistic and non-relativistic regimes (H. van Eerten, 2013; L. Nava et al., 2013). Afterglowpy supports various jet structures, including Gaussian and power-law jets, by decomposing the jet into multiple top-hat segments. The dynamical evolution of each segment is governed by coupled ordinary differential equations that describe the shock radius, dimensionless four-velocity, and evolving jet opening angle. These equations are integrated over time to generate synthetic afterglow light curves.
[…]
Parameter Prior Posterior
Top-Hat Gaussian Gaussian + WXT
$\log E_{K,\mathrm{iso}}$ (erg) (49, 55) $53.1^{+0.2}_{-0.2}$ $53.7^{+0.2}_{-0.2}$ $53.1^{+0.2}_{-0.1}$
$\log\theta_{c}$ (rad) (-3, -0.5) $0.8^{+0.1}_{-0.1}$ $1.2^{+0.2}_{-0.1}$ $0.9^{+0.02}_{-0.02}$
$\log\theta_{\rm obs}$ (rad) (-2, -0.5) <1.5<-1.5 $0.9^{+0.2}_{-0.2}$ 0.5-0.5 (Fixed)
$\log n$ (cm-3) (-6, 2) $3.7^{+0.7}_{-0.6}$ $2.6^{+1.1}_{-0.9}$ $0.3^{+0.3}_{-0.2}$
$p$ (2, 3.5) $2.42^{+0.03}_{-0.03}$ $2.38^{+0.04}_{-0.04}$ $2.32^{+0.07}_{-0.13}$
$\log\epsilon_{e}$ (-6, -0.5) $2^{+0.2}_{-0.1} 1.1$$1^{+0.2}_{-0.2}$ $0.62^{+0.08}_{-0.09}$
$\log\epsilon_{B}$ (-6, -0.5) >2.2>-2.2 $1.6^{+0.6}_{-0.8}$ $2.6^{+0.7}_{-0.7}$
$\log\Gamma_{0}$ (0.1, 5) $1.6^{+0.1}_{-0.1}$ >2.1>2.1 >2.5>2.5
$E_{K}$ ($10^{51}$ erg) $1.6^{+0.9}_{-0.6}$ $1.0^{+1.5}_{-0.4}$ $0.8^{+0.5}_{-0.3}$
We adopt a simple uniform (top-hat) jet profile to fit the afterglow data (see Fig.10) and treat the initial Lorentz factor $\Gamma_{0}$ as a free parameter [^4] A finite initial Lorentz factor breaks the assumption of instantaneous deceleration and necessitates disabling jet spreading, as standard spreading prescriptions are invalid under finite-$\Gamma_{0}$ conditions (G. Ryan et al., 2020). Accordingly, we set spread=False. For further discussion on the implications of finite initial Lorentz factors in Afterglowpy modeling, see R. Kaur et al. (2024).. The other free parameters include the isotropic-equivalent kinetic energy $E_{K,\mathrm{iso}}$, ambient density $n$, viewing angle $\theta_{\mathrm{obs}}$, jet core half-opening angle $\theta_{c}$, electron power-law index $p$, and the microphysical energy fractions in electrons ($\varepsilon_{e}$) and magnetic field ($\varepsilon_{B}$).
The posterior distributions of these parameters are summarized in Table 3 and visualized in Fig. 10. This model favors an on-axis jet ($\theta_{\rm obs}<\theta_{\rm c}$) with moderate initial Lorentz factor $\Gamma_{0}\sim$30-50. As our line of sight intercepts the jet core, we are not particularly sensitive to the jet structure, and the simple top-hat profile provides an adequate, although not unique, description. Adopting the median values of $E_{K,\mathrm{iso}}\sim 1.3\times 10^{53}\,\mathrm{erg}$ and the jet half opening angle $\theta_{c}\approx 0.16$ rad ($\sim 9^{\circ}$; see Fig. 10), the beaming corrected kinetic energy is $E_{K}=(1-\cos\theta_{c})E_{k,\mathrm{iso}}\approx 1.6\times 10^{51}\,\text{ erg}$, consistent with relativistic TDEs and typical GRB energy scales (X.-G. Wang et al., 2018; P. Beniamini et al., 2023).
We also tested the model of a structured jet with a Gaussian angular profile. This model provides an adequate description of the data, but points to a highly relativistic outflow with $\Gamma$>>130 seen off-axis ($\theta_{\rm obs}>\theta_{\rm c}$). Adopting the median values of $E_{K,\mathrm{iso}}\sim 5\times 10^{53}~\mathrm{erg}$ and a jet half-opening angle $\theta_{\rm c}\approx 0.06~\mathrm{rad}$ ($\sim 3.4^{\circ}$), the inferred beaming-corrected kinetic energy is $E_{K}\approx 10^{51}~\mathrm{erg}$. In this model, the jet core has a narrower opening angle than in the top-hat fit, but is surrounded by a wider angle structure from which emission is seen. Using this model, we can also reproduce the X-ray luminosity of the initial WXT detection (col. Gaussian+WXT). In this scenario, the early WXT flare would mark the afterglow onset rather than the time of the explosion.
In all cases, the derived blastwave kinetic energy is much higher than the upper limit on $E_{\gamma,\rm iso}\lesssim 8\times 10^{50}$ erg (Sect. 1). This results in a radiative efficiency which is unusually low for classical GRBs, $\eta_{\gamma}=E_{\gamma,\rm iso}/(E_{\gamma,\rm iso}+E_{\rm K,iso})\lesssim$1%, although sporadically observed in some events (e.g., GRB190829A, S. Dichiara et al. 2022; and the candidate off-axis GRB150101B, E. Troja et al. 2018)
6 Discussion
Our observations of EP241021a place it within a growing population of FXRTs that show compelling evidence for relativistic outflows despite the absence of prompt gamma-ray detections (e.g., S. B. Cenko et al., 2011; V. M. Lipunov et al., 2022; J. S. Bright et al., 2024; D. A. Perley et al., 2025; G. P. Srinivasaragavan et al., 2025; M. Busmann et al., 2025; W. X. Li et al., 2025). In Fig. 2 we show that most of these gamma-ray dark FXRTs require a soft photon index larger than 1 to explain the lack of gamma-rays, and identify a sub-class of events which implies unusually soft spectra with index $\gtrsim$1.5. EP241021a lies within the bulk of the FXRTs population, in a region that is also occupied by canonical GRBs. Therefore, the lack of gamma-ray signal does not immediately single it out as an outlier of the GRB population. Its moderate redshift ($z=0.75$) indicates that the lack of gamma-rays cannot be attributed to cosmological redshift effects, as might be the case for higher-redshift FXRTs (e.g., EP240315a at $z=4.9$; Y. Liu et al., 2025; R. Ricci et al., 2025). Instead, the emission of EP241021a is intrinsically soft, likely reflecting a low spectral peak or a suppressed gamma-ray output. The luminosity of its afterglow signal points to a highly energetic explosion with $E_{\rm K,iso}\gtrsim 10^{53}$ erg (Sect. 5), which requires an unusually low radiative efficiency $\eta_{\gamma}\lesssim 1\%$, at the lower end of the distribution found in typical GRBs (P. Beniamini & T. Piran, 2013; S. Dichiara et al., 2022).
The presence of both DISS and RISS in its radio afterglow reveals a compact emitting region, with angular size constraints of $\theta_{\text{src}}\gtrsim 1.5~\mu\mathrm{as}$ and a corresponding Lorentz factor of $\Gamma\approx 4$ at 1.5 d. Constraints coming from the brightness temperature (Sect. 4.3) and the optically thin radio spectrum (Sect. 5) provide independent support to this value. This evidence disfavors scenarios involving sub-relativistic ejecta, such as supernovae shock breakout and classical TDEs, and instead points to a relativistically expanding source, such as GRBs and relativistic TDEs. These observations demonstrate that EP241021a launched an outflow that is at least mildly relativistic, regardless of the lack of detectable gamma-ray emission. The absence of gamma rays thus does not preclude relativistic dynamics, and radio scintillation can serve as a powerful diagnostic for uncovering otherwise hidden jets.
Further support for this interpretation comes from our broadband afterglow modeling. The full radio data set, along with the early optical and X-ray data, can be reproduced by a standard forward shock model describing the interaction of a moderately relativistic jet with a low-density $n\approx 10^{-4}-10^{-1}~\mathrm{cm}^{-3}$ uniform medium. Two different scenarios are consistent with our dataset. The former scenario invokes a moderately relativistic ($\Gamma\approx$40) jet seen close to its axis (Top-Hat model in Table 3), in which the gamma-ray emission is likely suppressed by the compactness of the source (Y. Lithwick & R. Sari, 2001; T. Matsumoto et al., 2019). In this model, EP241021a differs from classical GRBs and resembles the properties of a dirty fireball (C. D. Dermer, 1999), where a baryon-rich outflow fails to produce bright $\gamma$-rays. This could point to a non-standard central engine, such as a magnetar or a relativistic TDE, launching an outflow with atypical properties.
The latter model invokes a highly relativistic structured jet seen off-axis (Gaussian model in Table 3). In this case, our line of sight is still dominated by the emission produced by ejecta expanding with moderate velocities. However, a high Lorentz factor ($\Gamma\gtrsim$130) jet exists and is not detected due to viewing angle effects. In this model, EP241021a could very well be a standard GRB “orphan” of its prompt emission (J. E. Rhoads, 1999). Within this off-axis framework, the origin of the initial WXT flare remains unclear. If our line of sight does not intercept the jet energetic core, then we might not be able to detect the traditional prompt emission phase. Thus, the luminous X-ray flare discovered by EP would be produced by the low-$\Gamma$ ejecta. Its onset could mark the break-out of the jet from the progenitor’s stellar envelope (C. D. Matzner, 2003; O. Bromberg et al., 2012), thus favoring the core-collapse of a massive star rather than a TDE. Alternatively, the WXT flare could represent the onset of the afterglow (S. Kobayashi & B. Zhang, 2007), produced as the relativistic ejecta begin to decelerate upon interacting with the ambient medium. The early X-ray signal would thus not require a separate emission component, but rather reflects the natural transition from the coasting to the deceleration phase in the jet’s evolution (Gaussian+WXT model, Table 3).
Our findings do not depend sensitively on the origin of the unusual optical rebrightening (M. Busmann et al., 2025), which substantially diverges from traditional afterglow models. Multiple interpretations were put forward to describe this atypical feature, such as refreshed shock (M. Busmann et al., 2025), multi-component jets (G. Gianfagna et al., 2025), or a reactivation of a long-lived central engine (X. Shu et al., 2025). Our analysis shows that the radio data are mostly consistent with a standard synchrotron radiation from a decelerating, moderately relativistic, wide-angled fireball, and are not substantially affected by any additional emission components probed at higher frequencies. To test the robustness of our results, we included the optical re-brightening and modeled the afterglow with a two-component Top-Hat jet, where both jets share the same $n$, $\theta_{c}$ and $\theta_{\mathrm{obs}}$ and have independent $E_{K,\mathrm{iso}}$, microphysical parameters and $\Gamma_{0}$. This shifts some of the inferred physical parameters (e.g., isotropic energy, circumburst density and viewing angle), but not the main inference of a wide-angled relativistic fireball (see Appendix).
Figure 8: Radio luminosity lightcurves. EP241021A (red circles) is compared to other FXRTs (e.g., EP240315A, R. Ricci et al. 2025; EP240414A, J. S. Bright et al. 2024) as well as other high-energy transients, including GRBs (circles; P. Chandra & D. A. Frail 2012), relativistic TDEs (squares; B. A. Zauderer et al. 2011; I. Andreoni et al. 2022) and FBOTs (diamonds; D. A. Perley et al. 2019; R. Margutti et al. 2019; D. A. Perley et al. 2021).
In Fig. 8 we place the radio counterpart of EP241021a in the broader context of high-energy transients. Its luminosity is substantially higher than FBOTs and low-luminosity GRBs (llGRBs), while the observed timescales differ from the population of merger-driven short GRBs. The radio light curve fits well within the distribution of cosmological GRBs from collapsars (dark grey circles) and relativistic jetted TDEs (squares). Its luminosity and timescale appear indeed similar to AT2022cmc (blue squares; I. Andreoni et al. 2022; D. R. Pasham et al. 2023).
Additional observations of gamma-ray dark FXRTs are essential to distinguish between different models and to understand the physical origin of this new population of high-energy transients. For instance, the total energetics, temporal profile and variability of the early X-ray emission detected by WXT would help clarify whether it is similar to erratic and highly variable prompt emission flares from an on-axis jet or consistent with other emission mechanisms visible from off-axis observers. Important clues might also come from the study of the environment, which would help differentiate between TDEs, core-collapse and other types of stellar explosions (e.g., A. S. Fruchter et al., 2006; S. van Velzen et al., 2021; B. O’Connor et al., 2025).
7 Conclusion
EP241021a is a luminous ($L_{X}\approx$ $9\times 10^{47}$ erg s-1) and relatively nearby ($z\approx$0.75) fast X-ray transient with no associated gamma-ray counterpart down to deep limits. We present long-term radio observations and afterglow modeling of this source, providing robust evidence for an outflow expanding with at least mildly relativistic velocities. Through a comprehensive monitoring campaign with ATCA and e-MERLIN over a period exceeding 100 days, we detect sustained radio emission and signatures of interstellar scintillation (ISS) at early times. This observed variability enables us to place constraints on the angular size and Lorentz factor of the emitting region, $\Gamma\approx$4 at 1.5 d.
Although we cannot rule out the presence of a highly relativistic jet oriented far from our line of sight, our modeling suggests that the observed emission comes from material moving toward us with moderately low initial Lorentz factor. This rules out progenitor models associated with sub-relativistic ejecta, and points to either relativistic TDEs or GRB-like explosions. Our findings reinforce the emerging view that some FXRTs may belong to a new population of energetic and at least moderately relativistic jets, lacking prompt gamma-ray emission due to geometric orientation or intrinsically low radiative inefficiency.
We thank David Williams for promptly scheduling, correlating and making the eMERLIN calibrated data available to the Authors. This work was supported by the European Research Council through the Consolidator grant BHianca (grant agreement ID 101002761). BO acknowledges useful discussions with Paz Beniamini, Ramandeep Gill, and Michael Moss. BO is supported by the McWilliams Fellowship at Carnegie Mellon University. YK is supported by the China Scholarship Council (CSC).
Appendix A Gamma-ray and X-ray flux values for Fast X-ray Transients
[…]
ID Name $\gamma$-ray flux Refγ X-ray flux RefX
($10^{-8}$ erg cm-2 s-1) ($10^{-11}$ erg cm-2 s-1)
1 EP240305a <3.8<3.8 This work $83_{-14}^{+20}$ M. J. Liu et al. 2024
2 EP240315a $118_{-36.8}^{+47.3}$ D. Svinkin et al. 2024b $53_{-7}^{+10}$ W. J. Zhang et al. 2024
3 EP240413a <2<2 This work $110\pm 10$ T. Y. Lian et al. 2024b
4 EP240417a <3.5<3.5 This work $9.8_{-8}^{+0.8}$ J. W. Hu et al. 2024d
5 EP240426b <4.0<4.0 This work $36\pm 6$ X. Pan et al. 2024a
6 EP240506a <6.2<6.2 This work $110_{-20}^{+30}$ D. Y. Li et al. 2024c
7 EP240617a <4.1<4.1 This work $350\pm 30$ H. Zhou et al. 2024d
8 EP240618a <0.85<0.85 M. E. Ravasio & P. G. Jonker 2024 $290_{-60}^{+70}$ H. Sun et al. 2024b
9 EP240625a <5.7<5.7 This work $29_{-5}^{+7}$ X. Pan et al. 2024b
10 EP240626a <4.8<4.8 This work $170_{-50}^{+73}$ Q. Y. Wu et al. 2024
11 EP240702a <4.3<4.3 This work $540_{-120}^{+150}$ W. Chen et al. 2024
12 EP240703a $39.4\pm 5.63$ D. Frederiks et al. 2024a $250_{-30}^{+40}$ Y. L. Wang et al. 2024a
13 EP240801a <3.6<3.6 This work $48\pm 31$ H. Zhou et al. 2024e
14 EP240802a $1249_{-40.0}^{+47.0}$ D. Frederiks et al. 2024b $170_{-30}^{+40}$ Y. L. Wang et al. 2024b
15 EP240804a $22\pm 9.45$ D. Frederiks et al. 2024c $61_{-16}^{+26}$ Y. Wang et al. 2024b
16 EP240807a $18.45_{-3.48}^{+2.32}$ A. Ridnaia et al. 2024 $17_{-6}^{+7}$ Y. C. Fu et al. 2024
17 EP240816a <9.3<9.3 This work $30\pm 5$ H. N. Yang et al. 2024
18 EP240820a <3.1<3.1 This work $12_{-5}^{+7}$ Y. F. Liang et al. 2024b
19 EP240913a $37\pm 0.2$ Fermi GBM Team 2024 $11_{-4}^{+5}$ D. Y. Li et al. 2024d
20 EP240918b <3.7<3.7 This work $26_{-12}^{+41}$ Y. F. Liang et al. 2024c
21 EP240918c <3.8<3.8 This work $150_{-20}^{+60}$ Y. F. Liang et al. 2024c
22 EP240919a $4.5\pm 0.5$ O. J. Roberts et al. 2024 $70_{-15}^{+18}$ Y. F. Liang et al. 2024d
23 EP240930a $26\pm 1$ J. Smith et al. 2024 $380\pm 50$ X. Tian et al. 2024
24 EP241021a <3.2<3.2 This work $33_{-16}^{+48}$ J. W. Hu et al. 2024c
25 EP241025a $11.8\pm 0.5$ M. Godwin & Fermi GBM Team 2024 $16\pm 7$ D. Y. Li et al. 2024b
26 EP241026a $41\pm 2$ A. Trigg & Fermi GBM Team 2024 $160\pm 30$ D. Y. Li et al. 2024a
27 EP241026b <4.8<4.8 This work $12_{-3}^{+4}$ T. Y. Lian et al. 2024a
28 EP241030a $37.6\pm 0.57$ C. de Barra et al. 2024 $7.5_{-2.4}^{+3}$ H. Z. Wu et al. 2024
29 EP241101a <3.2<3.2 This work $120_{-40}^{+50}$ Y. F. Liang et al. 2024a
30 EP241103a <5.5<5.5 This work $380_{-90}^{+110}$ T. Zhao et al. 2024
31 EP241104a $32\pm 7.6$ R. Sonawane et al. 2024 $20_{-11}^{+10}$ H. Zhou et al. 2024a
32 EP241107a <3.8<3.8 This work $10\pm 0.5$ R. Z. Li et al. 2024
33 EP241113a <4.0<4.0 This work $55_{-7.6}^{+12.6}$ Z. Y. Liu et al. 2024
34 EP241115a $15\pm 2$ D. Svinkin et al. (2024c) $14.2_{-4}^{+6}$ J. W. Hu et al. 2024a
35 EP241202b <4.5<4.5 This work $54_{-16}^{+20}$ C. Zhou et al. 2024
36 EP241206a <4.8<4.8 This work $49.2_{-12.4}^{+11.9}$ Y. H. I. Yin et al. 2024
37 EP241217a <4.1<4.1 This work $73\pm 27$ H. Zhou et al. 2024c
38 EP241217b $5.3\pm 0.2$ M. E. Ravasio et al. 2024 $119\pm 10$ H. Zhou et al. 2024b
39 EP250101a <8.0<8.0 This work $4.4_{-1.7}^{+2.3}$ Y. F. Liang et al. 2025
40 EP250108a <2.6<2.6 M. E. Ravasio et al. 2025 $4.2_{-0.9}^{+1.2}$ R. Z. Li et al. 2025a
41 EP250109a $13\pm 1$ U. Pathak et al. 2025a $250_{-120}^{+460}$ R. Z. Li et al. 2025b
42 EP250111a <8.0<8.0 This work $139_{-40}^{+53}$ T. Zhao et al. 2025
43 EP250125a <5.1<5.1 This work $180_{-50}^{+70}$ Q. Y. Wu et al. 2025
44 EP250205a $6.2\pm 0.2$ O. Mukherjee et al. 2025 $42\pm 11$ Z. Y. Liu et al. 2025
45 EP250207a <2.2<2.2 This work $29_{-7}^{+9}$ D. Y. Li et al. 2025b
46 EP250207b <5.4<5.4 This work $61_{-25}^{+42}$ X. Y. Zhou et al. 2025
47 EP250212a <3.8<3.8 This work $90_{-30}^{+100}$ D. Y. Li et al. 2025a
48 EP250223a <4.7<4.7 This work $44_{-11}^{+14}$ Y. Wang et al. 2025
49 EP250225a <3.7<3.7 This work $4.3_{-1.3}^{+2.1}$ S. Q. Jiang et al. 2025a
50 EP250226a $32\pm 1$ U. Pathak et al. 2025b $980$ S. Q. Jiang et al. 2025b
51 EP250227a <2.9<2.9 This work $210_{-120}^{+260}$ W. F. Wen et al. 2025
52 EP250302a <2.6<2.6 This work $700_{-160}^{+200}$ C. Y. Dai et al. 2025
53 EP250304a <3.9<3.9 This work $53\pm 4$ Y. J. Zhang et al. 2025
54 EP250321a <3.7<3.7 This work $173_{-19}^{+21}$ D. F. Hu et al. 2025
55 EP250321b <5.7<5.7 This work $7.6_{-3}^{+5.3}$ H. Q. Cheng et al. 2025
Appendix B Two-component TopHat Jet Fit (Including Optical Bump)
Figure 9: Broadband modeling including optical re-brightening results. Multi-band light curves fitted with a two-component uniform (Top-Hat) jet model. X-ray (Swift/XRT; 1 keV), optical/NIR ($r$, $z$, $J$; M. Busmann et al. 2025), and 15.5 GHz radio (AMI-LA; F. Carotenuto et al. 2024) data are shown. Downward arrows mark $3\sigma$ upper limits; open symbols indicate points excluded in the fit. Solid lines show the best-fitting two-component jet model, with contributions from component 1(dashed) and second component(dotted) indicated separately. Band-dependent scaling factors are applied for clarity.
Figure 10: Broadband modeling including optical re-brightening results. Posterior distributions of the fitted parameters for the uniform (Top-Hat) jet model.